OpenStudy (anonymous):

For the function f(x) = x^2 - 4x +5, x>= 2, which is equal to d/dx (f-1 this is the inverse)(x)) ? 1/(2y-4) where x and y are related by the equation (satisfy the equation) x=y^2-4y+5 x>= 1 2y-4 where x and y are related by the equation y= = x^2 - 4x +5, x>= 2 1/2x-4 for x>= 1 1/2x-4 for x>= 2 1/2y-4 where x and y are related by the equation y= = x^2 - 4x +5, x>= 2 inverse of x^2 - 4x +5 is 2±sqrt(x-1) and the derivative of that is (2x-4) but I don't know how to get the x>= 1 or x>= 2

OpenStudy (anonymous):

@amistre64 @mathmate

OpenStudy (anonymous):

@mathsciencehistory

OpenStudy (anonymous):

@NathanJHW

OpenStudy (phi):

There are 2 ways to decide between x>= 1 or x>= 2 FIrst, from your inverse function 2±sqrt(x-1) for a real number , you want x>=1 or, from the original function f(x) = x^2 - 4x +5, x>= 2 f(2) = 4-8+5= 1 so f(x) >= 1 and that means the inverse function's input (i.e. x ) >=1 btw, note that if we find dy/dx = 2x-4 then dx/dy is the derivative of the inverse function: 1/(2x-4)